Journalist Report – February 9th



Crew 274 Journalist Report 2-9-2023

Sol: 11

Author: Tony DiBernardo, Crew Journalist

Title: Take Me Home

Crew 274 woke up this morning at 7:30am to the theme from Apollo 13. The crew got their surveys done, then prepared for their last EVA of the mission.

EVA Activity

Today, HSO Pender retrieved his supply cache from the field and brough EEVAC-1 back home. He conducted performance metrics and final condition observations of the cache and overall, EEVAC-1 was found in good condition. The solar panels held up surprisingly well given the high winds from the previous evening. Finally, Pender and his EVA team collected the final data set on the orange radiation experiment led by Commander Sarah Guthrie.


Today, GHO Hines officially concluded both the primary and secondary growth experiments. While the secondary microgreen study proved to be a partial success, with the cress seed set resulting in a nearly 80% germination rate, only one seed among the several planted LDEF seeds were able to successfully germinate halfway through the mission. The resulting conclusion is being considered a success based on conservative expectations with plans to continue the study on future analogs.

Hab Design

Hab Specialist Bill O’Hara completed his study of MDRS architecture on schedule today. The result is a comprehensive set of notes regarding the awesome and inspiring design of the MDRS. He also supported the final EVA of the mission to bring home EEVAC-1.


Crew astronomers Loy and Pena observed detailed heliophysics phenomena and collected more bulk data sets on solar dark spots and solar chromosphere prominences. Large prominences were observed wrapping around solar magnetic fields, just before solar flares were registered by NOAA blasting away from a solar vortex. Eight large solar prominences and nine solar dark spots were observed today, which will be very useful for the astronomer’s analysis of solar cycle 25. On top of these observations, 60,000 images were captured, and 10,000 solar spots and prominences were stacked, measured, and analyzed.


59 separate 45-second long exposure images were collected from the MLCRCOS-16 New Mexico Observatory. These images are being separately captured through visual, infrared, and blue filters, assisting Astronomers Pena and Loy’s science on HADS variable star research. 59 B filter images were stacked, analyzed, and compared to previous images of this star system, HADS Variable Star V0799 AUR, to provide an assessment of this variable star’s brightness variability period and cycle, as well as provide a true-brightness standard to accurately measure distances in the immediate vicinity of HADS V0799 AUR’s stellar neighborhood.

EVA Report – February 9th



Crew 274 EVA Report 09-02-2023

EVA # 12

Author: Nicholas S. Pender, HSO

Purpose of EVA: Redeployment and data collection of EEVAC-1 (supply cache).

Start time: 10:00 AM

End time: 11:30 AM

Narrative: EVA #12’s purpose was to redeploy EEVAC-1 (the supply cache) located at coordinates N4248400, E519300 back to the MDRS habitat. Performance metrics collection and final condition observations of the cache were also conducted. EEVAC-1 was found in good condition. The solar panels held up surprisingly well given the high winds from the previous evening. We also collected the final data set on the orange radiation experiment led by Commander Sarah Guthrie. The EVA itself went very well, with no encounters with the public. One aspect that made this EVA different from others was the ingress process of the cache. Given the size of the cache, only two crew members could enter the airlock at a time. Noah and Salina ingressed first, carrying the solar panels and orange experiment materials. Nick and Tyler ingressed second carrying EEVAC-1. Listed below are the collected metrics for both experiments:

EEVAC-1 Metrics

  • Inside Temp (Thermostat): 67.4 F
  • Inside Temp (Data Logger): 62.7 F
  • Outside Temp: 57 F (likely due to position in sunlight)
  • Battery Level: 4 + 1 (charging)
  • Power Consumption: 0W
  • Total Consumption: 0.00 kWh

Orange Metrics

  • Inside Temp: 61 F
  • Outside Temp: 50 F
  • Inside radiation Lvl: 0 mSv
  • Outside radiation Lvl: 0 mSv

Destination: South on Cow Dung Road.

Coordinates (use UTM WGS 84): N4248400, E519300

Participants: Nicholas Pender (HSO), Noah Loy (Astronomer), Salina Pena (Astronomer), Tyler Hines (Greenhab)

Road(s) and routes per MDRS Map: EVA #12’s (map provided) purpose was to drive to EEVAC-1 deployed at N4248400, E519300 via south on Cow Dung Road near Zubrin’s Head.

Mode of travel: Rover: Spirit/Opportunity

Astronomy Report – February 9th



Astronomy Report

Name: Salina Peña and Noah Loy

Crew: 274

Date: 02/09/2023


Robotic Telescope Requested (Choose one): MDRS-14 MDRS-WF

Objects Viewed: Star/230209/variable star

Problems Encountered: None. Used the MLC RCOS 16 for observations.


Solar Features Observed: Sun/230209/solar prominence


Problems Encountered: None

GreenHab Report – February 9th



GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 80.425 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 14.79 gallons

Daily water usage for research and/or other purposes: 1.056 gallons

Water in Blue Tank: 44.044 gallons

Time(s) of watering for crops: 12:39 pm

Changes to crops: N/A

Narrative: Continued to actively maintain and water crops and plants as part of daily GreenHab operations. In addition, a culminating harvest of tomatoes was also performed to support the final phases of crew meal planning. Overall, general cleaning and trimming of the GreenHab were also conducted ahead of the crew cleaning the subsequent day.

Harvest: 86 grams of tomato

Support/supplies needed: N/A

Operations Report – February 9th



Crew 274 Operations Report 2/9/2023

SOL: 11

Name of person filing report: Alexis Joseph Lojek

Non-nominal systems: Toilet. Robotic Observatory.

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews. Robotic observatory is closed until further notice.


Spirit rover used: Yes

Hours: 208.8

Beginning charge: 100%

Ending charge: 53%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.9

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Curiosity rover used: No

Hours: 220.7

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: No

Hours: 255.4

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

General notes on rovers: Rovers Spirit and Opportunity were used for the morning EVA and plugged in when they returned, as it was a cloudless day and before noon. Curiosity and Perseverance were not utilized.

Summary of Hab operations: Toilet tank was emptied this evening and habitat is in a nominal condition.

WATER USE: 42 gallons as of 6:50pm

Water (static tank): 236 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged after EVA today and are off chargers. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 12:39 – 1:04 pm.

WATER USE: 14.79 gallons

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: Yes; 86 grams of tomatoes. Total: 86 grams.

Summary of ScienceDome operations: GHO Tyler Hines’ experiment concluded today. The Microgreens were a “partial success” according to Hines, with cress being the most successful to germinate. The LDEF seeds – one that was previously noted as germinating on Sol 6 is still germinating, and will be preserved for future study. Th rest of the seeds that did not germinate will be discarded.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The MDRS Robotic Observatory is inoperable until further notice due to the mechanical issue described in Sol 9’s Operations report.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Sol Summary Report – February 9th



Crew 274 Sol Summary Report 02-09-2023

Sol: 11

Summary Title: Finale

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: As the crew closes up research projects for this mission, they completed the last EVA this morning. Two teams retrieved the logistics cache which rested for a week in the “Martian” desert. The Greenhab Officer Tyler gathered a few more fresh tomatoes to add to the crew’s dinner and Crew Journalist DiBernardo wraps up final documentary videos and interviews of the crew. Astronomers Pena and Loy captured massive promenices and large sunspot groupings from the solar telescope bringing an exciting end to their projects.

Look Ahead Plan: The crew prepares for wrapping up its first mission at the Mars Desert Research Station with final reports, cleaning, and spending their last night together.

Anomalies in work: None

Weather: Sunny, 0C High, -8C Low; sunny

Crew Physical Status: Energetic

EVA: EVA #12, retrieval of the cache completed – successful

Reports to be filed: Sol Summary, Operations, EVA Report, GHO Report, and Astronomy Report.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site